OpenStudy (anonymous):

Find ln (-91.48) A. 4.5161 b. i pi - 4.5161 C. i pi +4.5161

OpenStudy (anonymous):

@amistre64 Sorry, if I'm tagging you too much, but no one is helping me on here besides you.

OpenStudy (anonymous):

log?

OpenStudy (anonymous):

oh i see a complex log use a calculator you will get C because it is always C

OpenStudy (anonymous):

http://www.wolframalpha.com/input/?i=log%28-91.48%29

OpenStudy (anonymous):

yes

jimthompson5910 (jim_thompson5910):

ln(-1) = pi*i where i = sqrt(-1) ----------------- ln(-91.48) = ln(-1*91.48) ln(-91.48) = ln(-1) + ln(91.48) ln(-91.48) = pi*i + 4.51612